OpenStudy (anonymous):

you are given a sample containing 7.5x10^6 cells/mL and told to prepare a sample of 3,000 cells/mL. What dilution factor is needed?

OpenStudy (anonymous):

2500

OpenStudy (xishem):

Using the dilution equation...\[C_1V_1=C_2V_2\]\[(7.5*10^6)V_1=(3*10^3)V_2\]\[V_2=2500(V_1) \] The volume of the first solution must be 2500 times less than the volume of the second solution, so the dilution factor is 2500.